OpenStudy (anonymous):

TRUE/FALSE There is no solution for csc x = 0.

OpenStudy (magbak):

What is the question please?

OpenStudy (anonymous):

True

OpenStudy (anonymous):

THAT'S CORRECT!!!

OpenStudy (anonymous):

Thats because \[\csc (\theta)= \frac{1}{\sin(\theta)}\]

OpenStudy (anonymous):

Can you divide 1 by any number to get 0?